OpenStudy (anonymous):

I GIVE MEDALS Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. A.) −3/4x + y = 10 B.) 3x − 4y = −40 C.) y = 3/4x + 12 D.) y = 3/4x + 10

OpenStudy (travisbrown372):

@turtleluver i love turtles too this is confusing :/

OpenStudy (anonymous):

...

OpenStudy (anonymous):

y-4=3/4(x+8) distribute the 2/4 to the x and 8 y-4=3/4x+6 add 4 y=3/4x+10

OpenStudy (anonymous):

are you sure that those choices are correct?

OpenStudy (anonymous):

lol yep its d

OpenStudy (anonymous):

srry...just forgot to type the fraction in that last question. XD

OpenStudy (anonymous):

ok